# Survey on Factuality in Large Language Models: Knowledge, Retrieval and Domain-Specificity

### v1

@@ 3.4 Evaluating Domain-specific Factuality
TABLE VI: Benchmarks for domain-specific factuality evaluation. The table presents the domain, tasks, datasets, and the LLMs evaluated in the respective study
Reference Domain Task Datasets Metrics Evaluated LLMs
Xie et al. (2023b) Finance
Sentiment analysis,
News headline classification
Named entity recognition
Question answering
Stock movement prediction
FLARE
F1, Acc,
Avg F1,
Entity F1,
EM, MCC
GPT-4 ,
BloombergGPT,
FinMA-(7B, 30B, 7B-full),
Vicuna-7B
Li et al. (2023e) Finance 134 E-com tasks
EcomInstruct
Micro-F1,
Macro-F1,
ROUGE
BLOOM, BLOOMZ,
ChatGPT, EcomGPT
Wang et al. (2023d) Medicine Multi-Choice QA CMB Acc
GPT-4, ChatGLM2-6B,
ChatGPT, DoctorGLM,
Baichuan-13B-chat,
HuatuoGPT, MedicalGPT,
ChatMed-Consult,
ChatGLM-Med ,
Bentsao, BianQue-2
Li et al. (2023b) Medicine Generative-QA Huatuo-26M
BLEU,
ROUGE,
GLEU
T5, GPT2
Jin et al. (2023) Medicine
Nomenclature,
Genomic location,
Functional analysis,
Sequence alignment
GeneTuring Acc
GPT-2, BioGPT,
BioMedLM, GPT-3,
ChatGPT, New Bing
Guha et al. (2023) Law
Issue-spotting,
Rule-recall,
Rule-application,
Rule-conclusion,
Interpretation,
Rhetorical-understanding
LegalBench
Acc, EM
GPT-4, GPT-3.5,
Claude-1, Incite, OPT
Falcon, LLaMA-2, FLAN-T5…
Fei et al. (2023) Law
Legal QA, NER,
Sentiment Analysis,
Reading Comprehension
LawBench
F1, Acc,
ROUGE-L,
Normalized
log-distance …
GPT-4,
ChatGPT,
InternLM-Chat,
StableBeluga2…
To assess the performance and factuality of these specialized LLMs, a plethora of datasets and benchmarks have been proposed across various domains. These resources not only serve as critical tools for evaluating the capabilities of LLMs but also facilitate advancements in specialized applications. We summarize them in Talbe VI. The distinction between this subsection and the previous two lies in its focus. This subsection delves deeper into factuality evaluation tailored to specific domains, while Sec 3.2 and 3.3 primarily concentrate on general factuality evaluation, with only a portion of their content dedicated to datasets evaluating factuality within specific domains.
@@
Xie et al. (2023b) designed a financial natural language understanding and prediction evaluation benchmark dubbed FLARE, based on their collected financial instruction tuning dataset FIT. This benchmark is used to evaluate their FinMA model. It randomly selects validation sets from FIT to choose the best model checkpoint and utilizes distinct test sets for evaluation. FLARE is a broader variant compared to the existing FLUE benchmark Shah et al. (2022) as it also encapsulates financial prediction tasks like stock movement prediction in addition to standard NLP tasks. The FLARE dataset includes several subtasks, such as sentiment analysis (FPB, FiQA-SA), news headline classification (Headline), named entity recognition (NER), question answering (FinQA, ConvFinQA), and stock movement prediction (BigData22, ACL18, CIKM18). Performance is gauged via a variety of metrics for each task, such as the accuracy and weighted F1 Score for sentiment analysis, entity-level F1 score for named entity recognition, and accuracy and Matthews correlation coefficient for stock movement prediction. Several methods, including their own FIT-fine-tuned FinMA and other LLMs (BloombergGPT, GPT-4, ChatGPT, BLOOM, GPT-NeoX, OPT-66B, Vicuna-13B) are dedicated to their comparison. BloombergGPT’s performance is assessed in various shot scenarios, while the zero-shot performance is reported for the remaining results. Some of the baselines depend on human evaluations given that LLMs without fine-tuning fail to generate instruction-defined answers. Conversely, FinMA’s results are conducted on a zero-shot basis and can be evaluated automatically. To enable direct comparison between the performance of FinMA and BloombergGPT, despite the former not releasing their test datasets, test datasets were constructed with the same data distribution.
Li et al. (2023e) propose EcomInstruct benchmark, which experiment investigates the performance of their EcomGPT language model in comparison to baseline models such as BLOOM and BLOOMZ. Categories of these baseline models include pre-trained large models with decoder-only architecture like BLOOM, and instruction-following language models like BLOOMZ and ChatGPT. The evaluation metric of the EcomInstruct involves converting all tasks to generative paradigms and using text generation evaluation metrics like ROUGE-L. Classification tasks were evaluated with precision, recall, and F1 scores. The EcomInstruct dataset, comprising 12 tasks across four major categories, is divided into training and testing sections. The EcomGPT is trained on 85,746 instances of E-commerce data. The performance of the model is assessed based on its capacity to generalize unseen tasks or datasets, with emphasis on cross-language and cross-task paradigm settings.
@@
Wang et al. (2023d) propose a localized medical benchmark called CMB, or the Comprehensive Medical Benchmark in Chinese. CMB is rooted entirely in the native Chinese linguistic and cultural framework. While traditional Chinese medicine is a significant part of this evaluation, it does not make up the entire benchmark. Both prominent LLMs, such as ChatGPT and GPT-4, and localized Chinese LLMs, including those specializing in the health domain, are evaluated using CMB. However, the benchmark is not designed as a leaderboard competition, but rather as a tool for self-assessment and understanding the progression of models in this field. CMB embodies a comprehensive, multi-layered medical benchmark in Chinese, comprised of hundreds of thousands of multiple-choice questions and complex case consultation questions. This wide range of queries covers all clinical medical specialties and various professional levels, seeking to evaluate a model’s medical knowledge and clinical consultation capabilities comprehensively.
@@
Li et al. (2023b) introduce Huatuo-26M dataset, the largest Chinese medical Question and Answer (QA) dataset to date, including over 26 million high-quality medical QA pairs. It covers a wide range of topics, such as diseases, symptoms, treatments, and drug information. The dataset is a valuable resource for anyone looking to improve AI applications in the medical field, such as chatbots and intelligent diagnostic systems. The Huatuo-26M dataset is gathered and integrated from various sources, including online medical encyclopedias, online medical knowledge bases, and online medical consultation records. Each QA pair in the dataset contains a problem description and a corresponding answer from a doctor or expert. Although a significant proportion of the Huatuo-26M dataset is constituted by the online medical consultation records, the text format data from these records is not publicly available, for unspecified reasons. This dataset is expected to be instrumental for multiple types of research and AI applications in the medical field. These areas of application extend to Natural Language Processing tasks like developing QA systems, text classification, sentiment analysis, and Machine Learning model training tasks like disease prediction and personalized treatment recommendation. Consequently, the dataset is favorable for developing AI applications in the medical field, from intelligent diagnosis systems to medical consultation chatbots.
@@
Jin et al. (2023) use nine tasks related to NCBI resources to evaluate the proposed GeneGPT model. The dataset used is GeneTuring benchmark Hou and Ji (2023), which contains 12 tasks with 50 question-answer pairs each. The tasks are divided into four modules - Nomenclature (Gene alias, Gene name conversion), Genomic location(Gene SNP association, Gene location, SNP location), Functional analysis(Gene disease association, Protein-coding genes), Sequence alignment (DNA to human genome, DNA to multiple species). Two settings of GeneGPT are assessed: a full setting where all prompt components are used, and a slim setting using only two components. The performance of GeneGPT is compared against various baselines including general-domain GPT-based LLMs like GPT-2, GPT-3, and ChatGPT. Additionally, GPT-2-sized biomedical domain-specific LLMs such as BioGPT and BioMedLM are evaluated. The new Bing, a retrieval-augmented LLM with access to relevant web pages, is also assessed. The result evaluation of the compared methods is based on the results reported in the original benchmark and are manually evaluated. However, the evaluation of the proposed GeneGPT method is determined through automatic evaluations.
@@
4 Analysis of Factuality
@@ 4.1 Analysis of Factuality
4.1.1 Knowledge Storage
The language model serves as a repository of knowledge, storing a multitude of information about the world within its parameters  Petroni et al. (2019b). However, the organization of this knowledge within LLMs remains largely mysterious. Meng et al. (2022) introduced a methodology called causal tracing to measure the indirect impact of hidden states or activations. This technique was employed to illustrate that factual knowledge is primarily stored in the early-layer feed-forward networks (FFNs) of such models. Similarly, Geva et al. (2021c) also suggests that a substantial portion of factual information is encoded within the FFN layers. They conceptualize the input of the FFN as a query, the first layer as keys, and the second layer as values. Consequently, the intermediate hidden dimension of the FFN can be interpreted as the number of memories within the layer, and the intermediate hidden state represents a vector comprising activation values for each memory. As a result, the final output of the FFN can be understood as the weighted sum of activated values. The authors further demonstrate that the value vectors often encapsulate human-interpretable concepts and knowledge Geva et al. (2022, 2023). In addition, Chen et al. (2023c) made an intriguing finding that the language model contains language-independent neurons that express multilingual knowledge and degenerate neurons that convey redundant information by applying the integrated gradients method Lundstrom et al. (2022). Nevertheless, it is important to note that the aforementioned studies primarily focus on the representation of individual facts, and the comprehensive understanding of how factual knowledge is precisely organized and interconnected within these models remains an ongoing challenge.
4.1.2 Knowledge Completeness and Awareness
@@
Several studies have investigated the knowledge awareness of Large Language Models, specifically assessing whether LLMs can accurately estimate the correctness of their own responses. Most of these studies treat LLMs as ”black boxes,” prompting the models to report their confidence levels or calculating the perplexity of the model’s output as an indicator of response likelihood. Gou et al. (2023) explore the model’s ability to validate and iteratively refine its outputs, akin to how humans interact with tools. The authors find that solely relying on self-correction without external feedback can lead to marginal improvements or even diminished performance. Ren et al. (2023) experiment with settings either augmented or not with external document retrieval to determine whether models recognize their own knowledge boundaries. Their findings indicate that LLMs possess an inaccurate perception of their factual knowledge boundaries and tend to be overly confident about their responses. LLMs often fail to fully harness the knowledge they possess; however, retrieval enhancement can somewhat compensate for this shortcoming. Yin et al. (2023) introduce a dataset named ”SelfAware” to test if models recognize what they don’t know, encompassing both answerable and unanswerable questions. The experiment suggests that models do possess some capacity to discern their own knowledge gaps, but they are still far from human levels. GPT-4 outperforms other models, instructions and In-Context-Learning Dong et al. (2023) can enhance a model’s discriminatory ability. Kadavath et al. (2022) focus on LLM self-assessment based on Language Model calibration using multiple-choice questions. Their findings revealed that the ”none of the above” option decreased accuracy, larger models showed better calibration, and RLHF hindered model calibration levels. However, simply adjusting the temperature parameter can rectify this issue. Azaria and Mitchell (2023) assess the truthfulness of statements generated by LLMs, by using the model’s internal state and hidden layer activations. The authors, employing a feedforward neural network, can classify if the model is misleading by utilizing the hidden output states.
@@
On the contrary, these observations have been contested in subsequent investigations. Kandpal et al. (2023) underscore the dependency of LLMs on the number of associated documents seen during pre-training. They argue that the success in answering fact-based questions is highly linked to the number of documents containing the topic of the question that were encountered in pre-training. The study further posited the necessity of scaling models extensively to achieve competitive performance for questions with minimum representation in the training data. Adding to these concerns, Sun et al. (2023b) critically evaluate the factual knowledge base of LLMs, using a specifically designed Head-to-Tail benchmark comprised of 18K question-answer pairs. The results show that the understanding of factual knowledge, particularly related to torso-to-tail entities, by currently available LLMs is suboptimal.
@@
4.1.3 Contextual Influence and Knowledge Conflict
@@
Some works explore the model’s capacity to utilize context, for example, Li et al. (2023a) observe that larger models tend to rely on their parametric knowledge, even when faced with counterfactual contexts. This suggests that as models increase in size, they might grow more confident in their internal knowledge, potentially sidelining external context. However, the introduction of irrelevant contexts can still influence their outputs. The balance between controllability (relying on relevant context) and robustness (resisting irrelevant context) emerges as a challenge in LLM training. The study indicates that reducing context noise improves controllability, but the effect on robustness remains to be seen. In contrast, Zhou et al. (2023b) propose prompt templates to guide LLMs towards more faithful text generation. Among these, opinion-based prompts prove most effective, indicating that when LLMs are queried about opinions, they adhere more closely to the context. Interestingly, the study finds that using counterfactual context enhances the model’s faithfulness, while original context, sourced from platforms like Wikipedia, might induce a simplicity bias, leading LLMs to answer questions without heavily relying on the context. Chen et al. (2023b) conduct a comprehensive evaluation of LLMs’ ability to effectively utilize retrieved information. The study reveals that while retrieved documents can boost LLM performance, the presence of noise in these documents can hinder it. Yue et al. (2023) investigate the nature of LLM-generated content in relation to provided references. They categorize the generated content as attributable, contradictory, or extrapolatory to the reference. Both fine-tuned models and instruction-based LLMs struggle to accurately evaluate the alignment between generated content and references, underscoring the challenge of ensuring that LLMs produce content consistent with the provided context.
@@
A series of studies are interested in LLMs’ behavior when confronted with conflicting information. Longpre et al. (2021) introduce the concept of knowledge conflicts, where the provided context contradicts the model’s learned information. Their findings suggest that such conflicts lead to increased prediction uncertainty, especially for in-domain examples. Observations across models, ranging from T5-60M to 11B, indicate that larger models tend to default to their parametric knowledge. Moreover, there’s an inverse relationship between retrieval quality and the tendency to rely on internal knowledge: the more irrelevant the evidence, the more the model defaults to its parametric knowledge. Chen et al. (2022) conduct experiments on typical ODQA models, including FiD and RAG. Their results show that FiD models rarely resort to memorization (less than 3.6% for NQ) compared to RAG models. Instead, FiD primarily grounds its answers in the provided evidence. Interestingly, when confronted with conflicting retrieved passages, models tend to fall back on their parametric knowledge. Xie et al. (2023a) explore the behavior of recent LLMs, including ChatGPT and GPT-4. Contrary to findings on smaller LMs, they discover that LLMs can be highly receptive to external evidence, even if it contradicts their parametric memory, provided the external evidence is coherent and convincing. Additionally, LLMs exhibit a strong confirmation bias, especially when presented with evidence that aligns with their parametric memory. This bias becomes even more pronounced for widely accepted knowledge. In scenarios where no relevant evidence is provided, LLMs tend to express uncertainty. However, when presented with both relevant and irrelevant evidence, they demonstrate an ability to filter out irrelevant information.
In conclusion, while studies like Li et al. (2023a) and Zhou et al. (2023b) emphasize the challenges and potential solutions in making LLMs more context-aware, others like Yue et al. (2023) and Xie et al. (2023a) highlight the inherent biases and limitations of LLMs. The overarching theme is the need for a balanced approach, where LLMs effectively leverage both their internal knowledge and external context to produce accurate and coherent outputs.
4.2 Causes of Factual Errors
Understanding the root causes of these factual inaccuracies is crucial for refining these models and ensuring their reliable application in real-world scenarios. In this subsection, we delve into the multifaceted origins of these errors, categorizing them based on the stages of model operation: Model Level, Retrieval Level, Generation Level, and other miscellaneous causes. Table I shows examples of factuality errors caused by different factors.
4.2.1 Model-level Causes

### v3

@@ 3.4. Evaluating Domain-specific Factuality
Table 7. Benchmarks for domain-specific factuality evaluation. The table presents the domain, tasks, datasets, and the LLMs evaluated in the respective study.
Reference Domain Task Metrics Evaluated LLMs FLARE (Xie et al., 2023a) Finance Sentiment analysis, News headline classification Named entity recognition Question answering Stock movement prediction F1, ACC, Avg F1, Entity F1, EM, MCC GPT-4 , BloombergGPT, FinMA-(7B, 30B, 7B-full), Vicuna-7B EcomInstruct (Li et al., 2023c) Finance 134 E-com tasks Micro-F1, Macro-F1, ROUGE BLOOM, BLOOMZ, ChatGPT, EcomGPT CMB (Wang et al., 2023a) Medicine Multi-Choice QA ACC GPT-4, ChatGLM2-6B, ChatGPT, DoctorGLM, Baichuan-13B-chat, HuatuoGPT, MedicalGPT, ChatMed-Consult, ChatGLM-Med , Bentsao, BianQue-2 Huatuo-26M (Li et al., 2023f) Medicine Generative-QA BLEU, ROUGE, GLEU T5, GPT2 NCBI (Jin et al., 2023) Medicine Nomenclature, Genomic location, Functional analysis, Sequence alignment ACC GPT-2, BioGPT, BioMedLM, GPT-3, ChatGPT, New Bing LegalBench (Guha et al., 2023) Law Issue-spotting, Rule-recall, Rule-application, Rule-conclusion, Interpretation, Rhetorical-understanding ACC, EM GPT-4, GPT-3.5, Claude-1, Incite, OPT Falcon, LLaMA-2, FLAN-T5… LawBench (Fei et al., 2023) Law Legal QA, NER, Sentiment Analysis, Reading Comprehension F1, ACC, ROUGE-L, Normalized log-distance … GPT-4, ChatGPT, InternLM-Chat, StableBeluga2… OceanBench (Bi et al., 2023) Geoscience 15 ocean- related tasks ACC GPT-4
In our examination of specialized Large Language Models (LLMs), we’ve identified an extensive array of datasets and benchmarks tailored to a variety of domains. These resources not only serve as critical tools for evaluating the capabilities of LLMs but also facilitate advancements in specialized applications. We summarize them in Table 7. The distinction between this subsection and the previous two lies in its focus. This subsection delves deeper into factuality evaluation tailored to specific domains, while Sec 3.2 and 3.3 primarily concentrate on general factuality evaluation, with only a portion of their content dedicated to datasets evaluating factuality within specific domains.
@@
Xie et al. (2023a) designed a financial natural language understanding and prediction evaluation benchmark dubbed FLARE, based on their collected financial instruction tuning dataset FIT. This benchmark is used to evaluate their FinMA model. It randomly selects validation sets from FIT to choose the best model checkpoint and utilizes distinct test sets for evaluation. FLARE is a broader variant compared to the existing FLUE benchmark (Shah et al., 2022) as it also encapsulates financial prediction tasks like stock movement prediction in addition to standard NLP tasks. The FLARE dataset includes several subtasks, such as sentiment analysis (FPB, FiQA-SA), news headline classification (Headline), named entity recognition (NER), question answering (FinQA, ConvFinQA), and stock movement prediction (BigData22, ACL18, CIKM18). Performance is gauged via a variety of metrics for each task, such as the accuracy and weighted F1 Score for sentiment analysis, entity-level F1 score for named entity recognition, and accuracy and Matthews correlation coefficient for stock movement prediction. Several methods, including their own FIT-fine-tuned FinMA and other LLMs (BloombergGPT, GPT-4, ChatGPT, BLOOM, GPT-NeoX, OPT-66B, Vicuna-13B) are dedicated to their comparison. BloombergGPT’s performance is assessed in various shot scenarios, while the zero-shot performance is reported for the remaining results. Some of the baselines depend on human evaluations given that LLMs without fine-tuning fail to generate instruction-defined answers. Conversely, FinMA’s results are conducted on a zero-shot basis and can be evaluated automatically. To enable direct comparison between the performance of FinMA and BloombergGPT, despite the former not releasing their test datasets, test datasets were constructed with the same data distribution.
Li et al. (2023c) propose EcomInstruct benchmark, which experiment investigates the performance of their EcomGPT language model in comparison to baseline models such as BLOOM and BLOOMZ. Categories of these baseline models include pre-trained large models with decoder-only architecture like BLOOM, and instruction-following language models like BLOOMZ and ChatGPT. The evaluation metric of the EcomInstruct involves converting all tasks to generative paradigms and using text generation evaluation metrics like ROUGE-L. Classification tasks were evaluated with precision, recall, and F1 scores. The EcomInstruct dataset, comprising 12 tasks across four major categories, is divided into training and testing sections. The EcomGPT is trained on 85,746 instances of E-commerce data. The performance of the model is assessed based on its capacity to generalize unseen tasks or datasets, with emphasis on cross-language and cross-task paradigm settings.
@@
Wang et al. (2023a) propose a localized medical benchmark called CMB, or the Comprehensive Medical Benchmark in Chinese. CMB is rooted entirely in the native Chinese linguistic and cultural framework. While traditional Chinese medicine is a significant part of this evaluation, it does not make up the entire benchmark. Both prominent LLMs, such as ChatGPT and GPT-4, and localized Chinese LLMs, including those specializing in the health domain, are evaluated using CMB. However, the benchmark is not designed as a leaderboard competition, but rather as a tool for self-assessment and understanding the progression of models in this field. CMB embodies a comprehensive, multi-layered medical benchmark in Chinese, comprised of hundreds of thousands of multiple-choice questions and complex case consultation questions. This wide range of queries covers all clinical medical specialties and various professional levels, seeking to evaluate a model’s medical knowledge and clinical consultation capabilities comprehensively.
@@
Li et al. (2023f) introduce Huatuo-26M dataset, the largest Chinese medical Question and Answer (QA) dataset to date, including over 26 million high-quality medical QA pairs. It covers a wide range of topics, such as diseases, symptoms, treatments, and drug information. The dataset is a valuable resource for anyone looking to improve AI applications in the medical field, such as chatbots and intelligent diagnostic systems. The Huatuo-26M dataset is gathered and integrated from various sources, including online medical encyclopedias, online medical knowledge bases, and online medical consultation records. Each QA pair in the dataset contains a problem description and a corresponding answer from a doctor or expert. Although a significant proportion of the Huatuo-26M dataset is constituted by the online medical consultation records, the text format data from these records is not publicly available, for unspecified reasons. This dataset is expected to be instrumental for multiple types of research and AI applications in the medical field. These areas of application extend to Natural Language Processing tasks like developing QA systems, text classification, sentiment analysis, and Machine Learning model training tasks like disease prediction and personalized treatment recommendation. Consequently, the dataset is favorable for developing AI applications in the medical field, from intelligent diagnosis systems to medical consultation chatbots.
@@
Jin et al. (2023) use nine tasks related to NCBI resources to evaluate the proposed GeneGPT model. The dataset used is GeneTuring benchmark (Hou and Ji, 2023), which contains 12 tasks with 50 question-answer pairs each. The tasks are divided into four modules - Nomenclature (Gene alias, Gene name conversion), Genomic location(Gene SNP association, Gene location, SNP location), Functional analysis(Gene disease association, Protein-coding genes), Sequence alignment (DNA to human genome, DNA to multiple species). Two settings of GeneGPT are assessed: a full setting where all prompt components are used, and a slim setting using only two components. The performance of GeneGPT is compared against various baselines including general-domain GPT-based LLMs like GPT-2, GPT-3, and ChatGPT. Additionally, GPT-2-sized biomedical domain-specific LLMs such as BioGPT and BioMedLM are evaluated. The new Bing, a retrieval-augmented LLM with access to relevant web pages, is also assessed. The result evaluation of the compared methods is based on the results reported in the original benchmark and are manually evaluated. However, the evaluation of the proposed GeneGPT method is determined through automatic evaluations.
@@
Bi et al. (2023) design OceanBench to evaluate the capabilities of LLMs for oceanography tasks. OceanBench includes a total of 15 type of ocean-related tasks, such as question-answering and description tasks. The samples in OceanBench are are automatically generated from the seed dataset and have undergone manual verification by experts.
4. Analysis of Factuality
@@ 4.1. Analysis of Factuality
4.1.1. Knowledge Storage
The language model serves as a repository of knowledge, storing a multitude of information about the world within its parameters  (Petroni et al., 2019b). However, the organization of this knowledge within LLMs remains largely mysterious. (Meng et al., 2022) introduce a methodology called causal tracing to measure the indirect impact of hidden states or activations. This technique was employed to illustrate that factual knowledge is primarily stored in the early-layer feed-forward networks (FFNs) of such models. Similarly, Geva et al. (2021c) also suggests that a substantial portion of factual information is encoded within the FFN layers. They conceptualize the input of the FFN as a query, the first layer as keys, and the second layer as values. Consequently, the intermediate hidden dimension of the FFN can be interpreted as the number of memories within the layer, and the intermediate hidden state represents a vector comprising activation values for each memory. As a result, the final output of the FFN can be understood as the weighted sum of activated values. The authors further demonstrate that the value vectors often encapsulate human-interpretable concepts and knowledge (Geva et al., 2022, 2023). In addition, Chen et al. (2023a) have made an intriguing finding that the language model contains language-independent neurons that express multilingual knowledge and degenerate neurons that convey redundant information by applying the integrated gradients method (Lundstrom et al., 2022). Nevertheless, it is important to note that the aforementioned studies primarily focus on the representation of individual facts, and the comprehensive understanding of how factual knowledge is precisely organized and interconnected within these models remains an ongoing challenge.
4.1.2. Knowledge Completeness and Awareness
@@
Several studies have investigated the knowledge awareness of Large Language Models, specifically assessing whether LLMs can accurately estimate the correctness of their own responses. Most of these studies treat LLMs as ”black boxes,” prompting the models to report their confidence levels or calculating the perplexity of the model’s output as an indicator of response likelihood. Gou et al. (2023) explore the model’s ability to validate and iteratively refine its outputs, akin to how humans interact with tools. The authors find that solely relying on self-correction without external feedback can lead to marginal improvements or even diminished performance. (Ren et al., 2023) experiment with settings either augmented or not with external document retrieval to determine whether models recognize their own knowledge boundaries. Their findings indicate that LLMs possess an inaccurate perception of their factual knowledge boundaries and tend to be overly confident about their responses. LLMs often fail to fully harness the knowledge they possess; however, retrieval enhancement can somewhat compensate for this shortcoming. Yin et al. (2023) introduce a dataset named “SelfAware” to test if models recognize what they don’t know, encompassing both answerable and unanswerable questions. The experiment suggests that models do possess some capacity to discern their own knowledge gaps, but they are still far from human levels. GPT-4 outperforms other models, instructions and In-Context-Learning (Dong et al., 2023) can enhance a model’s discriminatory ability. Kadavath et al. (2022) focus on LLM self-assessment based on Language Model calibration using multiple-choice questions. Their findings revealed that the ”none of the above” option decreased accuracy, larger models showed better calibration, and RLHF hindered model calibration levels. However, simply adjusting the temperature parameter can rectify this issue. Azaria and Mitchell (2023) assess the truthfulness of statements generated by LLMs, by using the model’s internal state and hidden layer activations. The authors, employing a feedforward neural network, can classify if the model is misleading by utilizing the hidden output states.
@@
On the contrary, these observations have been contested in subsequent investigations. Kandpal et al. (2023) underscore the dependency of LLMs on the number of associated documents seen during pre-training. They argue that the success in answering fact-based questions is highly linked to the number of documents containing the topic of the question that were encountered in pre-training. The study further posited the necessity of scaling models extensively to achieve competitive performance for questions with minimum representation in the training data. Adding to these concerns, Sun et al. (2023d) critically evaluate the factual knowledge base of LLMs, using a specifically designed Head-to-Tail benchmark comprised of 18K question-answer pairs. The results show that the understanding of factual knowledge, particularly related to torso-to-tail entities, by currently available LLMs is suboptimal.
@@
4.1.3. Contextual Influence and Knowledge Conflict
@@
Some works explore the model’s capacity to utilize context, for example, Li et al. (2023e) observe that larger models tend to rely on their parametric knowledge, even when faced with counterfactual contexts. This suggests that as models increase in size, they might grow more confident in their internal knowledge, potentially sidelining external context. However, the introduction of irrelevant contexts can still influence their outputs. The balance between controllability (relying on relevant context) and robustness (resisting irrelevant context) emerges as a challenge in LLM training. The study indicates that reducing context noise improves controllability, but the effect on robustness remains to be seen. In contrast, Zhou et al. (2023b) propose prompt templates to guide LLMs towards more faithful text generation. Among these, opinion-based prompts prove most effective, indicating that when LLMs are queried about opinions, they adhere more closely to the context. Interestingly, the study finds that using counterfactual context enhances the model’s faithfulness, while original context, sourced from platforms like Wikipedia, might induce a simplicity bias, leading LLMs to answer questions without heavily relying on the context. Chen et al. (2023d) conduct a comprehensive evaluation of LLMs’ ability to effectively utilize retrieved information. The study reveals that while retrieved documents can boost LLM performance, the presence of noise in these documents can hinder it. Yue et al. (2023) investigate the nature of LLM-generated content in relation to provided references. They categorize the generated content as attributable, contradictory, or extrapolatory to the reference. Both fine-tuned models and instruction-based LLMs struggle to accurately evaluate the alignment between generated content and references, underscoring the challenge of ensuring that LLMs produce content consistent with the provided context. Shi et al. (2023a) study the distractability of LLMs on the GSM-IC dataset derived from GSM8K (Cobbe et al., 2021). They discover that all the prompting techniques are responsive to irrelevant information in the problem definition. They identify various factors of irrelevant information that impact the model’s sensitivity to irrelevant context. Moreover, they find that self-consistency prompting and incorporating irrelevant information into the exemplars can enhance models’ performance, enabling them to learn to disregard irrelevant information.
@@
A series of studies are interested in LLMs’ behavior when confronted with conflicting information. Longpre et al. (2021) introduce the concept of knowledge conflicts, where the provided context contradicts the model’s learned information. Their findings suggest that such conflicts lead to increased prediction uncertainty, especially for in-domain examples. Observations across models, ranging from T5-60M to 11B, indicate that larger models tend to default to their parametric knowledge. Moreover, there’s an inverse relationship between retrieval quality and the tendency to rely on internal knowledge: the more irrelevant the evidence, the more the model defaults to its parametric knowledge. (Chen et al., 2022) conduct experiments on typical ODQA models, including FiD and RAG. Their results show that FiD models rarely resort to memorization (less than 3.6% for NQ) compared to RAG models. Instead, FiD primarily grounds its answers in the provided evidence. Interestingly, when confronted with conflicting retrieved passages, models tend to fall back on their parametric knowledge. (Xie et al., 2023b) explore the behavior of recent LLMs, including ChatGPT and GPT-4. Contrary to findings on smaller LMs, they discover that LLMs can be highly receptive to external evidence, even if it contradicts their parametric memory, provided the external evidence is coherent and convincing. Additionally, LLMs exhibit a strong confirmation bias, especially when presented with evidence that aligns with their parametric memory. This bias becomes even more pronounced for widely accepted knowledge. In scenarios where no relevant evidence is provided, LLMs tend to express uncertainty. However, when presented with both relevant and irrelevant evidence, they demonstrate an ability to filter out irrelevant information. Wang et al. (2023c) argue that LLMs should not rely solely on either parametric or non-parametric information, but grant LLM users the agency to make informed decisions. They introduce a framework including three tasks ((1) Contextual knowledge conflict detection; (2) QA-span knowledge conflict detection; (3) Distinct answers generation) to simulate knowledge conflicts and evaluate whether LLMs’ behaviors align with the goal.
In conclusion, while studies like Li et al. (2023e) and Zhou et al. (2023b) emphasize the challenges and potential solutions in making LLMs more context-aware, others like Yue et al. (2023) and Xie et al. (2023b) highlight the inherent biases and limitations of LLMs. The overarching theme is the need for a balanced approach, where LLMs effectively leverage both their internal knowledge and external context to produce accurate and coherent outputs.
4.2. Causes of Factual Errors
Understanding the root causes of these factual inaccuracies is crucial for refining these models and ensuring their reliable application in real-world scenarios. In this subsection, we delve into the multifaceted origins of these errors, categorizing them based on the stages of model operation: Model Level, Retrieval Level, Generation Level, and other miscellaneous causes. Table 1 shows examples of factuality errors caused by different factors.
4.2.1. Model-level Causes